OpenStudy (anonymous):

A juice can is 16cm tall and has a diameter of 9cm. The ends of the can are tin and the body is cardboard. A) what is the volume of the can? B) what area of tin is used? C) what area of cardboard is used? D) what is the total surface area?

OpenStudy (anonymous):

A) Volume= Area of base X Height=pi*radius^2 x height=(4.5)^2 x pi x 16=324pi

OpenStudy (anonymous):

Thanks alot! It would be nice if you can help me with more questions from my book! Thanks! I still have more!

OpenStudy (anonymous):

B)Since tin is used at both ends..and rest is cardboard's...so area of base=pi*radius^2=pi*(4.5)^2 area of tin used=2(area of base)=2(pi*(4.5)^2)

OpenStudy (anonymous):

How about c and d?

OpenStudy (anonymous):

c)surface area of sides=area of cardboard=2pi*radius*height find numeric value yourself

OpenStudy (anonymous):

d)total sufrace area=answer in part b+part c=area of sides + area of ends=2pi*radius*height+2(pi*radius^2)

OpenStudy (anonymous):

Thanks alot

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

Another question...a box of facial tissue is 22cm by 10.5 cm by 8cm. How much cardboard is on the outside surface?

OpenStudy (anonymous):

surface area of rectangular box=2lw + 2lh + 2wh which is the sum of areas of all six sides of the box..so plug in values

OpenStudy (anonymous):

Okay thanks you rock I have more if you don't mind

OpenStudy (anonymous):

you shouldn't be posting question within a question..make a new one

OpenStudy (anonymous):

A table is 60cm in diameter. A tablecloth hangs 65cm all around. A. What is the area of the tablecloth? B. how ,iChat lace trim is needed to edge the cloth? C. Lave trim is sold in whole meters at $2.89/m. How much would it cost to trim?

OpenStudy (anonymous):

Please just one more